OpenStudy (anonymous):

A sugar solution was made by mixing 7 ml of a 50% sugar solution and 3 ml of a 80% sugar solution. Find the concentration of the new mixture.

hero (hero):

amount of sugar equation: x + y = 10 percent of sugar equation: .50x + .80y = 10z We know x = 7, y = 3, therefore: 7 + 3 = 10 .50(7) + .80(3) = 10(z) Solve for z